# short-time Fourier transform iSTFT inverse short-time Fourier transform

### v1

@@ V-B4 Evaluation metrics
Objective metrics for evaluating singing voice restoration tasks are currently more limited compared to those available for speech processing. Following [34], we use the $\ell^{1}$-distance in the magnitude STFT domain and a Fréchet Audio Distance (FAD) using a VGGish embedding [86]. However, these are only limited in interpretability and hardly relate to listening impression [87, 88]. Therefore, we complete this evaluation benchmark with a listening test with 10 participants, using a similar setup as reported in Section V-A7. In this case, the test included 10 reverberant singing voice examples from the NHSS dataset, and the instructions were identical to those reported in Appendix -C1.
V-B5 Results
The results from the instrumental evaluation are reported in Table IV and those from the listening test in Fig. 4. The results show that BUDDy largely outperforms the unsupervised baselines and is on-par with the conditional diffusion model that benefited from supervision at training time. We found no statistically significant difference between the listening test scores of BUDDy and the conditional diffusion model (p-value of 0.55 in a paired Welch test).
@@ V-B5 Results
Figure 5: RIR estimation metrics calculated for each octave on the reverberant VCTK dataset. The violin plots show the results distribution, with the median highlighted. Lower values indicate better performance. FiNS [41] is trained in a supervision fashion whereas BUDDy is unsupervised.
V-C Room Impulse Response Estimation
@@
BUDDy is not only designed as a dereverberation algorithm but also functions as a blind unsupervised RIR estimator. We evaluate its performance for RIR estimation using the same speech model and data we employed for speech dereverberation in Section V-A.
V-C1 Baseline
@@
We benchmark BUDDy against FiNS [41], a deep neural network (DNN)-based approach trained to estimate time-domain RIRs directly from reverberant speech. FiNS comprises a 1D-convolutional encoder and a two-component decoder. The first decoder component estimates the late tail of the RIR by passing noise signals through a trainable filterbank containing several FIR filters. The second decoder component directly estimates the direct path and early reflections in the time-domain. In contrast to BUDDy, FiNS relies on supervised learning, thus requiring a paired dataset of reverberant speech and RIRs. We use an unofficial re-implementation [^2] https://github.com/kyungyunlee/fins and train the model on our VCTK-based reverberant speech dataset.
V-C2 Evaluation metrics
@@
Due to the highly ill-posed nature of the blind RIR estimation problem and the statistical nature of late reflections [11], we refrain from using element-wise distances, such as error-to-signal ratios, to evaluate the performance of RIR estimators. Instead, it is arguably more important to preserve the acoustic and perceptual properties of the reference RIR [89]. On the other hand, single metrics such as the T60 reverberation time or clarity index C50 do not account for frequency-specific estimation errors. We therefore employ a subband reverberation time T60 and clarity index C50, with subbands spanning octaves. This enables to keep a high-level representation of the acoustical properties while allowing enough granularity on the spectral attributes of the RIR. The reverberation time T60 is defined for a diffuse sound field as the time it takes for its energy decay curve (EDC) to decay by $60$dB [1]. To compute T60 from a RIR while avoiding the effects of the noise floor, we first determine T30 and extrapolate it to T60 by multiplying by a factor of 2. We calculate T30 as the time required for the EDC to decrease from -5 dB to -35 dB relative to the initial level to eliminate the influence of the direct path. This measure is computed in each octave band separately. The octave clarity index C50 is the ratio (in dB) between the energy in the $50$ first milliseconds and the energy in the remaining of the RIR, calculated in the corresponding octave band [1]. Consequently, we compute the absolute error between the T60 and C50 values calculated for each octave from the estimated RIR and those from the ground truth RIR.
V-C3 Results
@@
The results for both matched and mismatched test sets are plotted in Fig. 5. In the matched condition, FiNS and BUDDy achieve similar T60 error rates at low- and mid-range frequency bands, while BUDDy’s performance decreases at high frequencies (Fig. 5a). Our intuition is that the lower RIR estimation abilities of BUDDy at high frequencies can be linked to the tendency of diffusion models to generate high-frequency components at the later stages of the reverse diffusion process [90]. Consequently, there is less information available for optimizing the RIR parameters in this range, negatively affecting parameter convergence. A similar trend is observed for the C50 error (Fig. 5c). Furthermore, BUDDy generally achieves lower C50 error than FiNS in the mid-frequency range, where most of the speech content lies.
@@
In the mismatched setting, FiNS struggles to generalize because of its supervised training setup. As a result, BUDDy outperforms FiNS in both T60 and C50 error at low and mid-frequency bands (Figs. 5b) and 5d). At higher frequencies, BUDDy’s T60 estimation performance still remains slightly inferior to FiNS, though the gap is noticeably smaller than in the matched setting. Regarding C50, BUDDy outperforms FiNS on all frequency domains but the higher, 4-kHz-centered band. This increased relative performance of BUDDy compared to FiNS highlights the benefits of leveraging unsupervised training for RIR estimation in variable acoustic conditions.
VI Conclusion
@@
In this paper, we presented an unsupervised method that simultaneously performs blind dereverberation and RIR estimation using diffusion models. Our results highlight the importance of joint speech and RIR estimation in contrast to plugging estimated RIRs into informed dereverberation methods. The proposed method, BUDDy, yields state-of-the-art performance among unsupervised approaches for blind speech and singing voice dereverberation, outperforming both traditional and diffusion-based methods. Unlike blind supervised methods, which often struggle with generalization to unseen acoustic conditions, our unsupervised approach naturally overcomes this limitation due to its ability to adapt the reverberation operator to a broad range of room impulse responses. This holds as well for RIR estimation, as we show that the RIR estimation performance of BUDDy surpasses that of a state-of-the-art supervised DNN-based technique in mismatched acoustic conditions while being on par in a matched setting.
Acknowledgements
@@
We would like to thank Koichi Saito, Fumikuri Yohena and Kohei Yatabe for providing us with code and guidance through their methods.

### v2

@@ IV-B4 Evaluation metrics
Objective metrics for evaluating singing voice restoration tasks are limited compared to those available for speech processing. Following [34], we use the $\ell^{1}$-distance in the magnitude STFT domain and a Fréchet Audio Distance (FAD) using a VGGish embedding [90]. However, these are only limited in interpretability and hardly relate to listening impression [91, 92]. Therefore, we complete this evaluation benchmark with a listening test with 10 participants, using a similar setup as reported in Section IV-A7. The test included 12 reverberant singing voice examples from the NHSS dataset, containing 6 reverberant singing voice examples from each of the matched and mismatched datasets. The instructions were identical to those reported in Appendix -C1.
IV-B5 Results
The results from the instrumental evaluation are reported in Table IV and those from the listening test in Fig. 3. The results show that BUDDy largely outperforms the unsupervised baselines and the supervised baseline PSE on FAD and subjective listening. In particular, in the mismatched test set, PSE fails to produce high-quality speech whereas BUDDy remains consistent with its good performance on the matched test set. BUDDy outperforms SGMSE+ on FAD in the mismatched setting, However, although it loses a few points in the mismatched scenario, SGMSE+ surpasses BUDDy by a small yet statistically significant margin on both test benchmarks in the listening experiment.
IV-C Robustness of RIR-Informed Methods
Informed dereverberation algorithms such as [5, 26] assume complete knowledge of the room acoustics as provided by the RIR $\mathbf{h}$. However, as pointed out in Section I, even if the RIR is perfectly known, single channel dereverberation is not trivial as RIRs are mixed-phase systems, such that causal and stable inverse filters do not exist [12]. We examine here the sensitivity of informed dereverberation methods in partially blind scenarios, i.e. when RIRs are known up to estimation errors.
IV-C1 Baselines
We include several informed dereverberation approaches for comparison. The baseline InfDerevDPS is inspired from previous work [26] and fully described in Section II-B. We also include the regularized inverse filtering method RIF+Post [5]. The second method RIF+Post [5] performs regularized inverse filtering in the Fourier domain, followed by traditional speech enhancement [93]. More details on these baselines can be found in Appendix -C1.
IV-C2 Synthetic RIR estimation errors
We first study the case where the oracle RIR is corrupted by Gaussian noise. The results displayed in Fig. 4 indicate that the performance of both the diffusion-based and the traditional method dwindles as the noise power increases. This suggests that informed methods have very limited robustness to errors in the provided RIR. This is a crippling drawback since obtaining a perfect sample-wise estimation of RIRs is an arduous problem, given their stochastic nature [11].
IV-C3 DNN-based RIR estimation errors
We now shift to a more realistic scenario where the RIR is blindly estimated from the reverberant speech by a DNN, since in practice RIR estimation errors are unlikely to be perfectly Gaussian-distributed. In particular, we employ FiNS [41], a state-of-the-art supervised RIR estimator which obtains RIR estimates based on the reverberant utterance (see Section IV-D1 for details). We compare in Table V results where the RIR is perfectly known (i.e. informed scenario) versus when it is estimated by FiNS (i.e. partially blind). The acoustic conditions in the considered evaluation set match those of the training set. Therefore, since FiNS was trained in a supervised fashion using paired reverberant/RIR data, it is expected to perform well on such conditions. The dereverberation performance of both InfDerevDPS and RIF+Post is very poor when the RIR is estimated with FiNS [41], as opposed to when the RIR is perfectly known. Yet, through informal listening, we notice that FiNS produces perceptually reasonable RIR estimates, which highlights the very limited robustness of informed methods when estimation errors, even imperceptible, affect the RIR knowledge. This all suggests that in blind cases the RIR should be jointly estimated with the anechoic speech, which is the paradigm followed by our method BUDDy.
TABLE V: Dereverberation results on matched reverberant VCTK dataset. We indicate for each method in the table if it operates in a blind scenario.
Method Blind DNS-MOS PESQ
Reverberant - 3.14 $\pm$ 0.52 1.61 $\pm$ 0.37
RIF+Post [5] 3.41 $\pm$ 0.47 2.66 $\pm$ 0.40
InfDerevDPS [26] 3.91 $\pm$ 0.33 3.95 $\pm$ 0.42
FiNS/RIF+Post [41, 5] 2.18 $\pm$ 0.38 1.33 $\pm$ 0.19
FiNS/InfDerevDPS[41, 26] 2.19 $\pm$ 0.43 1.32 $\pm$ 0.18
BUDDy (ours) [40] 3.76 $\pm$ 0.41 2.30 $\pm$ 0.53
IV-D Room Impulse Response Estimation
@@
Figure 5: RIR estimation metrics for each octave and full-band on the reverberant VCTK dataset. The violin plots show the distribution and the median. Lower is better. FiNS [41] is trained in a supervision fashion whereas BUDDy is unsupervised.
@@
BUDDy is not only designed as a dereverberation algorithm but also functions as a blind unsupervised RIR estimator. We evaluate its performance for RIR estimation using the same speech model and data we employed for speech dereverberation in Section IV-A.
IV-D1 Baseline
@@
We benchmark BUDDy against FiNS [41], a DNN-based approach trained to estimate time-domain RIRs directly from reverberant speech. FiNS comprises a 1D-convolutional encoder and a two-component decoder. The first decoder component models the late tail of the RIR by passing noise signals through a trainable filterbank containing several FIR filters. The second decoder component directly estimates the direct path and early reflections in the time-domain. In contrast to BUDDy, FiNS relies on supervised learning, thus requiring a paired dataset of reverberant speech and RIRs. We use an unofficial re-implementation [^2] https://github.com/kyungyunlee/fins and train the model on our VCTK-based reverberant speech dataset.
IV-D2 Evaluation metrics
@@
Due to the highly ill-posed nature of the blind RIR estimation problem and the statistical nature of late reflections [11], we refrain from using element-wise distances, such as error-to-signal ratios, to evaluate the performance of RIR estimators. Instead, it is arguably more important to preserve the acoustic and perceptual properties of the reference RIR [94]. On the other hand, single metrics such as the full-band T60 reverberation time or clarity index C50 do not account for frequency-specific estimation errors. We therefore incorporate both full-band and subband reverberation time T60 and clarity index C50, with subbands spanning octaves. This enables to keep a high-level representation of the acoustical properties while allowing enough granularity on the spectral attributes of the RIR.
The reverberation time T60 is defined for a diffuse sound field as the time it takes for its energy decay curve (EDC) to decay by $60$ dB [1]. In order to avoid the effects of the noise floor, we calculate T60 as twice the time required for the EDC to decrease from $-5$ dB to $-35$ dB relative to the initial level, thereby eliminating the influence of the direct path. This measure is computed in each octave band separately. The octave clarity index C50 is the ratio (in dB) between the energy in the first $50$ ms and the energy in the remaining of the RIR, calculated in the corresponding octave band [1]. Consequently, we compute the absolute error between the T60 and C50 values calculated for each octave from the estimated RIR and those from the ground truth RIR.
IV-D3 Results
@@
The results for both matched and mismatched test sets are plotted in Fig. 5. In the matched condition, FiNS and BUDDy achieve similar T60 error rates at low- and mid-range frequency bands, while BUDDy’s performance decreases at high frequencies (Fig. 5a). Our intuition is that the lower RIR estimation abilities of BUDDy at high frequencies can be related to the tendency of diffusion models to generate high-frequency components in the later stages of the reverse diffusion process [95]. Consequently, there is less information available for optimizing the RIR parameters in this range when beginning sampling, negatively affecting parameter convergence. A similar trend is observed for the C50 error in Fig. 5c. Furthermore, BUDDy generally achieves lower C50 error than FiNS in the mid-frequency range, where most of the speech content lies.
@@
In the mismatched setting, FiNS struggles to generalize because of its supervised training setup. As a result, BUDDy outperforms FiNS in both T60 and C50 error at low and mid-frequency bands (Figs. 5b and 5d). At higher frequencies, BUDDy’s T60 estimation performance still remains slightly inferior to FiNS, though the gap is noticeably smaller than in the matched setting. Regarding C50, BUDDy outperforms FiNS in all frequency bands except for the highest 4-kHz band. This increased relative performance of BUDDy compared to FiNS highlights the benefits of leveraging unsupervised training for RIR estimation in variable acoustic conditions.
V Conclusion
@@
In this paper, we presented an unsupervised method that simultaneously performs blind dereverberation and RIR estimation using diffusion models. Our results highlight the importance of joint speech and RIR estimation in contrast to plugging estimated RIRs into informed dereverberation methods. The proposed method, BUDDy, yields state-of-the-art performance among unsupervised approaches for blind speech and singing voice dereverberation, outperforming both traditional and DNN-based methods. Unlike blind supervised methods, which often struggle with generalization to unseen acoustic conditions, our unsupervised approach naturally overcomes this limitation due to its ability to adapt the reverberation operator to a broad range of RIRs. This holds as well for RIR estimation, as we show that the RIR estimation performance of BUDDy surpasses that of a state-of-the-art supervised DNN-based technique in mismatched acoustic conditions while being on par in a matched setting.
Acknowledgments
@@
We would like to thank Koichi Saito, Fumikuri Yohena, and Kohei Yatabe for providing us with code and guidance through their methods. Many thanks to Julius Richter and Till Svajda for their advice on retraining SGMSE+ for singing voice dereverberation.